OpenStudy (anonymous):

Suppose that an object moves along an s-axis so that its location is given by s(t)=t^2+3t at time t. (Here s is in meters and t is in seconds.) Find the average velocity of the object in meters per second over the time interval t=1 to t=9 seconds. (in m/s) Find the instantaneous velocity of the object in meters per second at t=4 seconds. (in m/s)

OpenStudy (dumbcow):

v(t) = s'(t) s'(t) = 2t + 3 v(4) = 2(4) +3 = 11 instantaneous velocity at t=4 is 11m/s avg velocity is total distance/total time avg velocity = s(9)-s(1)/(9-1) s(9) = 9^2 +(9)(3) = 81+27 = 108 s(1) = 1+3 = 4 avg velocity = (108-4)/8 = 104/8 = 13 m/s

OpenStudy (anonymous):

urgh, stupid errors again on my part. You're the man.

OpenStudy (dumbcow):

yeah watch out for the little mistakes, but they happen to everyone keep it up :)